OpenStudy (anonymous):

If a buffer solution is .110 M in a weak acid (Ka = 2.7*10^-5) and .530 M in its conjugate base, what is the pH?

OpenStudy (anonymous):

\[pH = p Ka + \log \left[ A-/ HA \right]\] \[pKA = - \log KA = - \log (2.7\times10^{5}) = 5.431\] \[pH =5.431 + \log \left( 0.110/0.530\right)\] \[pH = 5.431 -0.68 = 4.751\]